# On multi-graded Proj schemes

### v4

@@ 2.3 Potions
    If $Q$ is a graded $A$-module, we will also denote by $Q_{(S)}$ the degree-00 part $(Q_{S})_{0}$ of $Q_{S}$; it admits a canonical structure of $A_{(S)}$-module.
@@
    We have a canonical morphism of potion rings $A_{(S)}\to A_{(ST)}$.
@@
    Assume that $S$ is relevant, and that $T$ does not contain 00. Fix a subset $T^{\prime}\subset T$ which generates $T$ as a submonoid of $(A,\times)$ and, for any $t$ in $T^{\prime}$, fix $n_{t}\in\mathbb{N}_{>0}$ and $s_{t},s_{t}^{\prime}\in\underline{S}$ nonzero such that $\deg(t^{n_{t}})=\deg(s_{t})-\deg(s_{t}^{\prime})$. Then $\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}$ belongs to the potion $A_{(\underline{S})}=A_{(S)}$. Moreover we have a canonical isomorphism of $A_{(S)}$-algebras between $A_{(ST)}$ and the localization of $A_{(S)}$ with respect to the multiplicative subset of $A_{(S)}$ generated by the elements $\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}$.
@@
    $\mathrm{Spec}(A_{(ST)})\to\mathrm{Spec}(A_{(S)})$
    induced by the ring morphism in (1) is an open immersion of schemes.
@@
    ${\operatorname{Spec}}(A_{(f_{1}+\cdots+f_{n})})\to{\operatorname{Spec}}(A_{(f_{1})})\cup\cdots\cup{\operatorname{Spec}}(A_{(f_{n})})$
@@
(1) The desired morphism is obtained as the degree-zero part of the canonical morphism of graded rings $A_{S}\to A_{ST}$.
(2) Without loss of generality, we can (and will) assume that $S=\underline{S}$. Consider the localization $\big{(}A_{(S)}\big{)}_{\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}}$ of $A_{(S)}$ with respect to the homogeneous multiplicative subset generated by $\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}$. Since the image of $\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}$ in $A_{(ST)}$ is invertible for any $t\in T^{\prime}$, the universal property of localizations gives us a canonical morphism of $A_{(S)}$-algebras
@@
where $(A_{(f_{1}+\cdots+f_{n})})_{f_{i}^{\prime}}$ is the localization of $A_{(f_{1}+\cdots+f_{n})}$ with respect to the multiplicative subset generated by $f_{i}^{\prime}$. We have $f_{1}^{\prime}+\cdots+f_{n}^{\prime}=1$ in $A_{(f_{1}+\cdots+f_{n})}$, so by Iten(3) in [StP, Tag 01HS] we obtain that
@@ 2.4 Examples
Let $A$ be a commutative ring and let $\{[M_{i},a_{i}]:i\in I\}$ be a multi-center in $A$ in the sense of [Ma23b], i.e. $M_{i}$ is an ideal of $A$ and $a_{i}$ is an element of $A$ for any $i\in I$. For $i\in I$, let $L_{i}$ be the ideal $M_{i}+(a_{i})$ of $A$. Let ${\operatorname{Bl}}_{\{L_{i}:i\in I\}}A=\bigoplus_{\nu\in{\mathbb{N}}_{I}}L^{\nu}$ be the multi-Rees $A$-algebra associated with $A$ and $\{L_{i}:i\in I\}$. (Here, if $\nu=(\nu_{i}:i\in I)$, then $L^{\nu}$ is the product of the $(L_{i})^{\nu_{i}}$’s.) The ring ${\operatorname{Bl}}_{\{L_{i}:i\in I\}}A$ is naturally ${\mathbb{Z}}_{I}$-graded, and each $a_{i}$ defines an element in its degree-$e_{i}$ part, also denoted $a_{i}$. Then the family $\{a_{i}:i\in I\}$ is relevant, and [Ma23b, Fact 2.29] shows that the potion ring $({\operatorname{Bl}}_{\{L_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}$ identifies with the dilatation $A[\{\frac{M_{i}}{a_{i}}:i\in I\}]$.
@@ 3.1 Proj scheme of a graded ring
In [BS03, Definition 2.2], Brenner and Schröer define the Proj scheme [^3] Brenner–Schröer use the terminology “homogeneous spectrum” of $A$. We prefer to use the term “Proj scheme” to emphasize the parallel with Grothendieck’s construction. associated with $A$ as a certain open subset in the ringed space $\mathrm{Quot}(A)$ obtained as the cokernel (in the sense of [SGA3, Exp. V, §1.b]) of the maps $a$ and $p$. In Construction 3.1 below we explain an equivalent description of this scheme, obtained by gluing spectra of certain potion rings. (See §3.2 for a justification that our description is indeed equivalent to that in [BS03].)
@@
    As explained in Remark 2.4(1), if the subgroup of $M$ generated by degrees of nonzero homogeneous elements in $A$ is of finite index, $\{0\}\in{\mathcal{F}}_{A}$. However, the localization $A_{\{0\}}$ is the zero ring, so that $D_{\dagger}(\{0\})$ is empty.
@@ 3.2 First properties
By construction, ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $D_{\dagger}(S)$, which have the property that the intersection of any two such subschemes is affine. This implies that the diagonal morphism of ${\operatorname{Proj}}^{M}(A)$ is affine (as in [BS03, Proposition 3.1]), hence (since affine morphisms are quasi-compact) the following statement.
@@
For any $S\in\mathcal{F}_{A}$ we have a canonical morphism $A_{0}\to A_{(S)}$; these morphisms glue to a morphism of schemes
${\operatorname{Proj}}^{M}(A)\to{\operatorname{Spec}}(A_{0}).$ (3.1)
@@
    In the $\mathbb{N}$-graded setting, the scheme ${\operatorname{Proj}}^{\mathbb{Z}}(A)$ is always separated; see [StP, Tag 01MC]. For general $M$ and $A$ this is not true; for separatedness criteria, see [BS03, §3] and Proposition 3.7 below.
@@
${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Proj}}^{M}(A).$ (3.4)
@@
$H\times_{{\operatorname{Spec}}(A_{0})}\bigl{(}{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\bigr{)}\to{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\to{\operatorname{Proj}}^{M}(A)$
factors through a morphism $H\times_{{\operatorname{Spec}}(A_{0})}{\operatorname{Proj}}^{M}(A)\to{\operatorname{Proj}}^{M}(A)$. It is easily seen that this morphism defines an action of $H$ on ${\operatorname{Proj}}^{M}(A)$.
@@ 3.3 Open subschemes defined by quasi-relevant elements
${\operatorname{Spec}}(A_{(a)})\to\bigcup_{i=1}^{n}D_{\dagger}(f_{i}).$
@@
${\operatorname{Spec}}(A_{(a)})\to{\operatorname{Proj}}^{M}(A).$ (3.5)
@@
${\operatorname{Spec}}(A_{(a)})=\bigcup_{i=1}^{n}{\operatorname{Spec}}(A_{(af_{i})}),\quad\text{resp.}\quad{\operatorname{Spec}}(A_{(a)})=\bigcup_{j=1}^{m}{\operatorname{Spec}}(A_{(ag_{i})})$
@@ 3.4 More basic properties
We start with a functoriality property of the Proj scheme construction, which generalizes [StP, Tag 01MY]. For this we note that if $A,B$ are commutative $M$-graded rings and $S\subset A$ is a homogeneous multiplicative subset, then so is $\Psi(S)$. Clearly, $\Psi(S)$ is relevant if $S$ is, and finitely generated as a submonoid of $(B,\times)$ if $S$ is finitely generated as a submonoid of $(A,\times)$. Hence $\Psi$ induces a map $\mathcal{F}_{A}\to\mathcal{F}_{B}$, which will also be denoted $\Psi$.
@@
Let $\Psi:A\to B$ be a morphism of $M$-graded rings. For any $\mathcal{F}\subset\mathcal{F}_{A}$ we have a canonical morphism of schemes $\mathrm{Proj}^{M}_{\Psi(\mathcal{F})}(B)\to\mathrm{Proj}^{M}_{\mathcal{F}}(A)$. Moreover, for any $S\in{\mathcal{F}}$ we have
@@
The morphism is obtained by glueing from the morphisms induced by the canonical morphisms $A_{(S)}\to B_{(\Psi(S))}$ for $S\in\mathcal{F}$.
@@
is a closed immersion. This follows from (3.6), since the morphism $A_{(f)}\to B_{(\Psi(f))}$ is clearly surjective. ∎
@@
Let $R$ be a commutative ring and $N$ be a commutative cancellative monoid such that the group $M=N^{\mathrm{gp}}$ is finitely generated. (See §1.6.3 for the terminology and notation.) Let $R[N]:=\bigoplus_{m\in N}R\cdot X^{m}$ be the $R$-algebra of the monoid $N$, with its natural $M$-grading. The homogeneous elements in this ring are those of the form $rX^{m}$ with $r\in R$ and $m\in N$; such an element is relevant if and only if $N\cap(m-N)$ generates $M$. For each such element we have $(R[N])_{(rX^{m})}=R_{r}$; therefore we have $\mathrm{Proj}(R[N])=\mathrm{Spec}(R)$.

### v5

@@ 2.3 Potions
    If $Q$ is a graded $A$-module, we will also denote by $Q_{(S)}$ the degree-00 part $(Q_{S})_{0}$ of $Q_{S}$; it admits a canonical structure of an $A_{(S)}$-module.
@@
    We have a canonical homomorphism of potion rings $A_{(S)}\to A_{(ST)}$.
@@
    Assume that $S$ is relevant. Fix a subset $T^{\prime}\subset T$ which generates $T$ as a submonoid of $(A,\times)$ and, for any $t$ in $T^{\prime}$, fix $n_{t}\in\mathbb{N}_{>0}$ and $s_{t},s_{t}^{\prime}\in\underline{S}$ such that $\deg(t^{n_{t}})=\deg(s_{t})-\deg(s_{t}^{\prime})$. Then $\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}$ belongs to the potion $A_{(\underline{S})}=A_{(S)}$. Moreover we have a canonical isomorphism of $A_{(S)}$-algebras between $A_{(ST)}$ and the localization of $A_{(S)}$ with respect to the multiplicative subset of $A_{(S)}$ generated by the elements $\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}$.
@@
    $\mathrm{Spec}(A_{(ST)})\longrightarrow\mathrm{Spec}(A_{(S)})$
    induced by the ring homomorphism in (1) is an open immersion of schemes.
@@
    ${\operatorname{Spec}}(A_{(f_{1}+\cdots+f_{n})})\longrightarrow{\operatorname{Spec}}(A_{(f_{1})})\cup\cdots\cup{\operatorname{Spec}}(A_{(f_{n})})$
@@
(1) The desired morphism is obtained as the degree-zero part of the canonical homomorphism of graded rings $A_{S}\to A_{ST}$.
(2) Without loss of generality, we can (and will) assume that $S=\underline{S}$. Consider the localization $\big{(}A_{(S)}\big{)}_{\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}}$ of $A_{(S)}$ with respect to the homogeneous multiplicative subset generated by $\{\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}:t\in T^{\prime}\}$. Since the image of $\frac{t^{n_{t}}s_{t}^{\prime}}{s_{t}}$ in $A_{(ST)}$ is invertible for any $t\in T^{\prime}$, the universal property of localizations gives us a canonical homomorphism of $A_{(S)}$-algebras
@@
where $(A_{(f_{1}+\cdots+f_{n})})_{f_{i}^{\prime}}$ is the localization of $A_{(f_{1}+\cdots+f_{n})}$ with respect to the multiplicative subset generated by $f_{i}^{\prime}$. We have $f_{1}^{\prime}+\cdots+f_{n}^{\prime}=1$ in $A_{(f_{1}+\cdots+f_{n})}$, so by (3) in [StP, Tag 01HS] we obtain that
@@ 2.4 Examples
Let $A$ be a commutative ring and let $\{[M_{i},a_{i}]:i\in I\}$ be a multi-center in $A$ in the sense of [Ma24], i.e. $M_{i}$ is an ideal of $A$ and $a_{i}$ is an element of $A$ for any $i\in I$. For $i\in I$, let $L_{i}$ be the ideal $M_{i}+(a_{i})$ of $A$. Let ${\operatorname{Bl}}_{\{L_{i}:i\in I\}}A=\bigoplus_{\nu\in{\mathbb{N}}_{I}}L^{\nu}$ be the multi-Rees $A$-algebra associated with $A$ and $\{L_{i}:i\in I\}$. (Here, if $\nu=(\nu_{i}:i\in I)$, then $L^{\nu}$ is the product of the $(L_{i})^{\nu_{i}}$’s.) The ring ${\operatorname{Bl}}_{\{L_{i}:i\in I\}}A$ is naturally ${\mathbb{Z}}_{I}$-graded, and each $a_{i}$ defines an element in its degree-$e_{i}$ part, also denoted $a_{i}$. Then the family $\{a_{i}:i\in I\}$ is relevant, and [Ma24, Fact 2.35] shows that the potion ring $({\operatorname{Bl}}_{\{L_{i}:i\in I\}}A)_{(\{a_{i}:i\in I\})}$ identifies with the dilatation $A[\{\frac{M_{i}}{a_{i}}:i\in I\}]$.
@@ 3.1 Proj scheme of a graded ring
In [BS03, Definition 2.2], Brenner and Schröer define the Proj scheme [^3] Brenner–Schröer use the terminology “homogeneous spectrum” of $A$. We prefer to use the term “Proj scheme” to emphasize the parallel with Grothendieck’s construction. associated with $A$ as a certain open subset in the ringed space $\mathrm{Quot}(A)$ obtained as the cokernel (in the sense of [SGA3, Exp. V, §1.b]) of the maps $a$ and $p$. In Construction 3.1 below, we explain an equivalent description of this scheme, obtained by gluing spectra of certain potion rings. (See §3.2 for a justification that our description is indeed equivalent to that in [BS03].)
@@
    As explained in Remark 2.4, $\{0\}\in{\mathcal{F}}_{A}$. However, the localization $A_{\{0\}}$ is the zero ring, so that $D_{\dagger}(\{0\})$ is empty.
@@
  4. 4.
    Assume that $\{1\}$ is relevant. Then $D_{\dagger}(\{1\})={\operatorname{Spec}}(A_{0})$ by definition. Moreover, for any relevant family $S$, we have $S=S\{1\}$ and $D_{\dagger}(S)=\left(D_{\dagger}(S)\cap D_{\dagger}(\{1\})\right)\subset D_{\dagger}(\{1\})$. So ${\operatorname{Proj}}^{M}(A)={\operatorname{Spec}}(A_{0})$.
@@
Remark 3.4.
If $A$ is any commutative ring, let $\mathrm{Prim}(A)$ be the set of prime ideals of $A$, without topology. For any subset $T\subset A$, let $D(T)\subset\mathrm{Prim}(A)$ be the set $\{\mathfrak{p}\in\mathrm{Prim}(A):\forall t\in T,\,t\not\in\mathfrak{p}\}.$ Let $\mathrm{Alex}(A)$ be the set $\mathrm{Prim}(A)$, endowed with the topology generated by all the $D(T)$ for $T\subset A$. Then $\mathrm{Alex}(A)$ is an Alexandroff topological space and in fact the Alexandroffication of ${\operatorname{Spec}}(A)$. We call the topology of $\mathrm{Alex}(A)$ the Alexandroff–Zariski topology. In general, this is not the discrete topology, nor the Zariski topology (e.g. take $A={\mathbb{Z}}$).
Recall that the association $D(f)\mapsto A_{f}$ (for $f\in A$) extends to a sheaf of rings on ${\operatorname{Spec}}(A)$, namely the structure sheaf ${\mathcal{O}}_{{\operatorname{Spec}}(A)}$ [StP, Tag 01HU]. It is a natural and elementary question to ask whether the association $D(T)\mapsto A_{T}$ (for $T\subset A$) extends to a sheaf of rings on $\mathrm{Alex}(A)$, refining ${\mathcal{O}}_{{\operatorname{Spec}}(A)}$. In other words, considering the continuous Alexandroffication map $f:\mathrm{Alex}(A)\to{\operatorname{Spec}}(A)$, does the pullback presheaf $f_{p}{\mathcal{O}}_{{\operatorname{Spec}}(A)}$ coincide with the pullback sheaf $f^{-1}{\mathcal{O}}_{{\operatorname{Spec}}(A)}$? (See e.g. [StP, Tag 008C] for the notation $f_{p}$ and $f^{-1}$.) Even if it is easy to see that the answer is Yes in many cases (e.g. if $A$ is integral), the answer is negative in general. For example take $A=({\mathbb{Z}}/2{\mathbb{Z}})^{\mathbb{N}}$ (Boolean ring). In this case $\mathrm{Alex}(A)$ is the discrete topology, so that while the global sections of $f_{p}{\mathcal{O}}_{{\operatorname{Spec}}(A)}$ give $A$, the global sections of $f^{-1}{\mathcal{O}}_{{\operatorname{Spec}}(A)}$ are products of ${\mathbb{Z}}/2{\mathbb{Z}}$ over $\mathrm{Prim}(A)$. It is well-known that $\mathrm{Prim}(A)$ is in bijection with the Stone–Cech compactification of ${\mathbb{N}}$, which has ordinality strictly bigger than ${\mathbb{N}}$; as a consequence, $f_{p}{\mathcal{O}}_{{\operatorname{Spec}}(A)}\neq f^{-1}{\mathcal{O}}_{{\operatorname{Spec}}(A)}$.
Therefore, one cannot refine the structure sheaf of affine schemes (and the theory of schemes) in this way, using arbitrary localizations. So it is not possible to create a theory, similar to the theory of schemes in the framework of locally ringed space, allowing glueing along open subspaces in the Alexandroff–Zariski topology, which justifies Remark 1.2. Note that Exercise 11.1 of [Sch19] says that one can produce a theory inspired by the theory of schemes using arbitrary localizations, but it is not a theory based on locally ringed spaces, and will not behave formally as scheme theory regarding glueings. Note that the theory of Analytic Spaces of [Sch19] uses sorts of quasi-compact coverings, so that one cannot define the Proj of an arbirary graded ring in this theory either. In the above geometric paradigms, it thus seems that the hypothesis that $A$ is graded by a finitely generated abelian group cannot be removed without serious difficulties to define ${\operatorname{Proj}}$ schemes.
@@ 3.2 First properties
By construction, ${\operatorname{Proj}}^{M}(A)$ is covered by the affine open subschemes $D_{\dagger}(S)$, which have the property that the intersection of any two such subschemes is affine. This implies that the diagonal morphism of ${\operatorname{Proj}}^{M}(A)$ is affine (as in [BS03, Proposition 3.1]), hence (since affine morphisms are quasi-compact) we get the following statement.
@@
For any $S\in\mathcal{F}_{A}$ we have a canonical homomorphism $A_{0}\to A_{(S)}$; these homomorphisms glue to a morphism of schemes
${\operatorname{Proj}}^{M}(A)\longrightarrow{\operatorname{Spec}}(A_{0}).$ (3.1)
@@
    In the $\mathbb{N}$-graded setting, the scheme ${\operatorname{Proj}}^{\mathbb{Z}}(A)$ is always separated; see [StP, Tag 01MC]. For general $M$ and $A$ this is not true, even when $M={\mathbb{Z}}$. For examples and separatedness criteria, see [BS03, §3] and Proposition 3.8 below.
@@
${\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\longrightarrow{\operatorname{Proj}}^{M}(A).$ (3.4)
@@
$H\times_{{\operatorname{Spec}}(A_{0})}\bigl{(}{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\bigr{)}\longrightarrow{\operatorname{Spec}}(A)\smallsetminus V(A_{\dagger})\longrightarrow{\operatorname{Proj}}^{M}(A)$
factors through a morphism $H\times_{{\operatorname{Spec}}(A_{0})}{\operatorname{Proj}}^{M}(A)\longrightarrow{\operatorname{Proj}}^{M}(A)$. It is easily seen that this morphism defines an action of $H$ on ${\operatorname{Proj}}^{M}(A)$.
@@ 3.3 Open subschemes defined by quasi-relevant elements
${\operatorname{Spec}}(A_{(a)})\longrightarrow\bigcup_{i=1}^{n}D_{\dagger}(f_{i}).$
@@
${\operatorname{Spec}}(A_{(a)})\longrightarrow{\operatorname{Proj}}^{M}(A).$ (3.5)
@@
${\operatorname{Spec}}(A_{(a)})=\bigcup_{i=1}^{n}{\operatorname{Spec}}(A_{(af_{i})}),\quad\text{resp.}\quad{\operatorname{Spec}}(A_{(a)})=\bigcup_{j=1}^{m}{\operatorname{Spec}}(A_{(ag_{j})})$
@@ 3.4 More basic properties
We start with a functoriality property of the Proj scheme construction, which generalizes [StP, Tag 01MY]. For this we note that if $\Psi:A\to B$ is a homomorphism of commutative $M$-graded rings and $S\subset A$ is a homogeneous multiplicative subset, then so is $\Psi(S)$. Clearly, $\Psi(S)$ is relevant if $S$ is, and finitely generated as a submonoid of $(B,\times)$ if $S$ is finitely generated as a submonoid of $(A,\times)$. Hence $\Psi$ induces a map $\mathcal{F}_{A}\to\mathcal{F}_{B}$, which will also be denoted $\Psi$.
@@
Let $\Psi:A\to B$ be a homomorphism of $M$-graded rings. For any $\mathcal{F}\subset\mathcal{F}_{A}$ we have a canonical morphism of schemes $\mathrm{Proj}^{M}_{\Psi(\mathcal{F})}(B)\to\mathrm{Proj}^{M}_{\mathcal{F}}(A)$. Moreover, for any $S\in{\mathcal{F}}$ we have
@@
The morphism is obtained by glueing from the morphisms induced by the canonical homomorphisms $A_{(S)}\to B_{(\Psi(S))}$ for $S\in\mathcal{F}$.
@@
is a closed immersion. This follows from (3.6), since the homomorphism $A_{(f)}\to B_{(\Psi(f))}$ is clearly surjective. ∎
@@
Example 3.17.
Let $A={\mathbb{Z}}[X_{1},\ldots,X_{n+1}]$ be ${\mathbb{Z}}$-graded by $\deg(X_{1}^{k_{1}}\cdots X_{n+1}^{k_{n+1}})=\sum_{j=1}^{n+1}k_{j}$. Let $A^{\prime}={\mathbb{Z}}[X_{1},\ldots,X_{n^{\prime}+1}]$ be similarly ${\mathbb{Z}}$-graded. We have ${\operatorname{Proj}}^{\mathbb{Z}}(A)={\mathbb{P}}^{n}$ and ${\operatorname{Proj}}^{\mathbb{Z}}(A^{\prime})={\mathbb{P}}^{n^{\prime}}$. Proposition 3.16 implies that ${\mathbb{P}}^{n}\times_{{\operatorname{Spec}}({\mathbb{Z}})}{\mathbb{P}}^{n^{\prime}}={\operatorname{Proj}}^{{\mathbb{Z}}^{2}}(A\otimes_{{\mathbb{Z}}}A^{\prime})$
@@
Let $R$ be a commutative ring and $N$ be a commutative cancellative monoid such that the group $M=N^{\mathrm{gp}}$ is finitely generated. Let $R[N]:=\bigoplus_{m\in N}R\cdot X^{m}$ be the $R$-algebra of the monoid $N$, with its natural $M$-grading. The homogeneous elements in this ring are those of the form $rX^{m}$ with $r\in R$ and $m\in N$; such an element is relevant if and only if $N\cap(m-N)$ generates $M$ up to torsion. For each such element we have $(R[N])_{(rX^{m})}=R_{r}$; therefore we have $\mathrm{Proj}(R[N])=\mathrm{Spec}(R)$.
It is a standard fact that, even in the $\mathbb{N}$-graded setting, very different graded rings can have the same Proj scheme, see e.g. the discussion of the Veronese embedding in [EG]. We generalize this observation in the following proposition.